# GRaD-Nav: Efficiently Learning Visual Drone Navigation with Gaussian Radiance Fields and Differentiable Dynamics

### v2

@@ Abstract
Autonomous visual navigation is an essential element in robot autonomy. Reinforcement learning (RL) offers a promising policy training paradigm. However existing RL methods suffer from high sample complexity, poor sim-to-real transfer, and limited runtime adaptability to navigation scenarios not seen during training. These problems are particularly challenging for drones, with complex nonlinear and unstable dynamics, and strong dynamic coupling between control and perception. In this paper, we propose a novel framework that integrates 3D Gaussian Splatting (3DGS) with differentiable deep reinforcement learning (DDRL) to train vision-based drone navigation policies. By leveraging high-fidelity 3D scene representations and differentiable simulation, our method improves sample efficiency and sim-to-real transfer. Additionally, we incorporate a Context-aided Estimator Network (CENet) to adapt to environmental variations at runtime. Moreover, by curriculum training in a mixture of different surrounding environments, we achieve in-task generalization, the ability to solve new instances of a task not seen during training. Drone hardware experiments demonstrate our method’s high training efficiency compared to state-of-the-art RL methods, zero shot sim-to-real transfer for real robot deployment without fine tuning, and ability to adapt to new instances within the same task class (e.g. to fly through a gate at different locations with different distractors in the environment).
@@ I Introduction
Autonomous drones have the potential to improve practices in agriculture, environmental management, and search and rescue [1, 2, 3]. A critical aspect of these robots’ functionality is their efficient navigation and control in complex, unstructured environments [4]. Traditional approaches to this problem have predominantly relied on a stack of different modules including perception, localization, mapping, planning, and control [5], [6]. However, the integration of these different modules has many issues, including high system complexity and computational overhead, communication latency between modules, multiple points of failure, and difficult-to-characterize error propagation between modules.
@@
  • We introduce a simulator for training robot visuomotor policies by integrating 3DGS for high-fidelity visuals with a differentiable dynamics model for computing end-to-end gradients.
@@ II-A2 Short-Horizon Actor-Critic
where $\tilde{V}(\mathbf{s})$ is the estimated value of state $\mathrm{s}$, and is computed from the sampled short-horizon trajectories through a td-$\lambda$ formulation [32].
@@ III-A3 Implementation datails
We used a standard open-source codebase, Nerfstudio [33] as our 3DGS training and inference platform. The differentiable drone dynamics model is also implemented with PyTorch, which enables efficient Jacobian computation through autograd for training the policy using our GRaD-Nav algorithm. When parallel simulating 128 drones flying in a highly unstructured and cluttered area (room size $\approx 100\,m^{2}$, 3DGS model size $\approx$ 1.5GB), setting the simulated time step as 0.05s (20Hz), the simulation wall-clock time is about 0.7s per step, tested on a desktop with i9-13900K CPU, RTX 4090 GPU, and 48GB RAM. The detailed account of each item’s time consumption during simulation is listed in Table I.
@@ III-C1 Reward function
Linear velocity $\|\bm{v}\|^{2}$ $-0.5$
Pose $\|\bm{q}-\bm{q}_{0}\|$ $-0.5$
Height $(\bm{h}-\bm{h}_{\text{target}})^{2}$ $-2.0$
Action $\|\bm{a}_{t}\|^{2}$ $-1.0$
Action rate $\|\bm{a}_{t}-\bm{a}_{t-1}\|^{2}$ $-1.0$
Smoothness $\|\bm{a}_{t}-2\bm{a}_{t-1}+\bm{a}_{t-2}\|^{2}$ $-1.0$
@@
[…]
Waypoint $\exp(-\bm{d}_{\text{wp}})$ $2.0$
Obstacle avoidance $\bm{d}_{\text{obst}}$ $1.0$
Out-of-map $\bm{x}_{\text{o.b.}}^{2}+\bm{y}_{\text{o.b.}}^{2}$ $-2.0$
Ref. traj. tracking $\|\frac{\bm{v}}{\|\bm{v}\|}-\frac{\bm{v}_{\text{des}}}{\|\bm{v}_{\text{ref}}\|}\|$ $-2.0$
@@ III-C2 Domain randomization
Mass $[1.0,1.5]$ kg
Max. thrust $[22,30]$ N
$I_{x},I_{y}$ inertial $[0.075,0.125]$ $\text{kg}\cdot m^{2}$
$I_{z}$ inertial $[0.15,0.25]$ $\text{kg}\cdot m^{2}$
Motor delay factor $[0.5,0.8]$
Body rate delay factor $[0.5,0.8]$
Drag force coefficient $[0.4,0.6]$
@@ III-C3 Guiding drone to fly desired trajectories
Additionally, to ensure safe navigation, we incorporate an obstacle avoidance reward $r_{\text{obs}}$, which encourages the drone to maintain a safe distance from obstacles visible in its FOV. The reward formulation is as follows:
$r_{\text{obs}}=\bm{d}_{\text{obs}},\quad\bm{d}_{\text{obs}}<\bm{d}_{\text{th}},$ (17)
where $\bm{d}_{\text{obs}}$ represents the minimum distance to the nearest obstacle in drone’s FOV, $\bm{d}_{\text{th}}$ is the predefined threshold for proximity, which is set to 0.5m in our method.
@@ IV-A Training efficiency comparison with other methods
To validate GRaD-Nav’s training efficiency, we compared our method with other differentiable and non-differentiable RL algorithms. Back Propagation Through Time (BPTT) [36] is a critic-free differentiable reinforcement learning algorithm that passes gradients through the whole trajectory; Proximal Policy Optimization (PPO) [37] is a common non-differentiable, online reinforcement learning algorithm. All three algorithms were used for training policies that navigate through a trajectory of around 20m, passing through several challenging unstructured obstacles, which can be seen in 3. As much as possible, we used identical hyperparameters (Table VI) for different algorithms. It is to be noted that BPTT samples the whole trajectory for policy updating, meaning the horizon length equals the episode length, which can take up a lot of memory for each environment and heavily limits its maximum allowed number of parallel simulated environments. The experiment results in Fig.2 show that (i) non-differentiable RL can struggle to train a satisfactory policy for this end-to-end visual navigation task; (ii) compared to vanilla BPTT, our method achieves over 300% sample efficiency and uses only 20% of training time to deliver a better policy.
@@ IV-B Ablation study of our methods
Trajectories Middle gate Right gate
@@
The experiment results show that our proposed method achieves the highest training and evaluation rewards as well as success rate on both trajectories among all methods. As visual perception is our navigation policy’s major sensor input, it is not surprising that the policy without visual observation cannot conduct successful navigation. Using depth image rather than RGB image also demonstrated a low success rate in vision-ablated experiments, which emphasized the necessity of accessing RGB information for effective navigation on our task. As introduced in Section III-C3, drone’s body linear velocity plays an important role in our policy’s navigation-related reward items, velocity-ablated policies would suffer from serious compounding error, and thus cannot fly an ideal trajectory. Without CENet, our method can still train a policy network that achieves high rewards compared to other ablation cases. CENet ablated policies can complete a trajectory successfully when the initial condition is close to the ”neutral” initial condition (e.g. position = (0,0,1.3)m). However, policies without CENet can be vulnerable to any environmental perturbations, which seriously harm the robustness of the policy networks. All of the failure cases without CENet on two trajectories ”crash” due to unsuccessful obstacle avoidance.
@@ IV-C Sim-to-real transfer of generalizable policy
Figure 4: Generalizable policy to fly through gates at different positions and with different distractor objects.
@@ V Conclusions
Limitations: Our method requires significant reward shaping, e.g. with hand-engineered trajectory waypoints. This also leads to policies that are adept at a single task (e.g., flying through a gate), but not able to execute multiple tasks. Our policy also requires a reliable onboard estimate of velocity, which can be obtained from lower-level VIO modules, which can give inconsistent performance in practice. Future work can focus on various aspects, including (i) training a multi-task policy with language conditioning; (ii) further improving our policy’s generalization ability by upgrading backbone policy network and training in a greater varieties of surrounding environments; (iii) extending this work to more contact-rich scenarios like holistic mobile manipulation and aerial manipulation.
@@ VI Acknowledgment
The authors would like to thank Chenghao Zhu for assistance with the PPO baseline code development. The authors would also like to thank Keiko Nagami and Javier Yu for fruitful discussion on FiGS and sharing with NeRF usage code. The GRaD-Nav algorithm was developed based on SHAC’s open-sourced codebase [29], $\beta$-VAE based CENet code was referred to [38].

### v3

@@ Abstract
Autonomous visual navigation is an essential element in robot autonomy. Reinforcement learning (RL) offers a promising policy training paradigm. However, existing RL methods suffer from high sample complexity, poor sim-to-real transfer, and limited runtime adaptability. These problems are particularly challenging for drones, with complex nonlinear and unstable dynamics, and strong dynamic coupling between control and perception. In this paper, we propose a novel framework that integrates 3D Gaussian Splatting (3DGS) with differentiable deep reinforcement learning (DDRL) to train vision-based drone navigation policies. By leveraging high-fidelity 3D scene representations and differentiable simulation, our method improves sample efficiency and sim-to-real transfer. Additionally, we incorporate a Context-aided Estimator Network (CENet) to adapt to environmental variations at runtime. Moreover, by curriculum training in a mixture of different surrounding environments, we achieve in-task generalization, the ability to solve new instances of a task not seen during training. Drone hardware experiments demonstrate our method’s high training efficiency compared to state-of-the-art RL methods, zero shot sim-to-real transfer for real robot deployment without fine tuning, and ability to adapt to new instances within the same task class (e.g. to fly through a gate at different locations with different distractors in the environment). Our simulator and training framework are open-sourced at: https://github.com/Qianzhong-Chen/grad_nav.
@@ I Introduction
Autonomous drones have the potential to improve practices in agriculture, environmental management, and search and rescue [1, 2, 3]. A critical aspect of these robots’ functionality is their efficient navigation and control in complex, unstructured environments [4]. Traditional approaches to this problem have predominantly relied on a stack of different modules including perception, localization, mapping, planning, and control [5, 6, 7]. However, the integration of these different modules has many issues, including high system complexity and computational overhead, communication latency between modules, multiple points of failure, and difficult-to-characterize error propagation between modules.
@@
  • We introduce a simulator for training robot vision-based control policies by integrating 3DGS for high-fidelity visuals with a differentiable dynamics model to enable end-to-end gradient computation.
@@ II-A2 Short-Horizon Actor-Critic
[…]
@@ III-A3 Implementation datails
We used a standard open-source codebase, Nerfstudio [34] as our 3DGS training and inference platform. The differentiable drone dynamics model is also implemented with PyTorch, which enables efficient Jacobian computation through autograd for training the policy using our GRaD-Nav algorithm. When parallel simulating 128 drones flying in a highly unstructured and cluttered area (room size $\approx 100\,m^{2}$, 3DGS model size $\approx$ 1.5GB), setting the simulated time step as 0.05s (20Hz), the simulation wall-clock time is about 0.07s per step, tested on a desktop with i9-13900K CPU, RTX 4090 GPU, and 48GB RAM. The detailed account of each item’s time consumption during simulation is listed in Table I.
@@ III-C1 Reward function
Linear velocity $\|\bm{v}\|^{2}$ 0.5-0.5- 0.5
Pose $\|\bm{q}-\bm{q}_{0}\|$ 0.5-0.5- 0.5
Height $(\bm{h}-\bm{h}_{\text{target}})^{2}$ 2.0-2.0- 2.0
Action $\|\bm{a}_{t}\|^{2}$ 1.0-1.0- 1.0
Action rate $\|\bm{a}_{t}-\bm{a}_{t-1}\|^{2}$ 1.0-1.0- 1.0
Smoothness $\|\bm{a}_{t}-2\bm{a}_{t-1}+\bm{a}_{t-2}\|^{2}$ 1.0-1.0- 1.0
@@
Yaw alignment $\frac{\bm{v}_{xy}}{\|\bm{v}_{xy}\|}\cdot\hat{\bm{y}}_{\text{yaw}}$ 0.250.250.25
Waypoint $\exp(-\bm{d}_{\text{wp}})$ 2.02.02.0
Obstacle avoidance $\bm{d}_{\text{obst}}$ 1.01.01.0
Out-of-map $\bm{x}_{\text{o.b.}}^{2}+\bm{y}_{\text{o.b.}}^{2}$ 2.0-2.0- 2.0
Ref. traj. tracking $\|\frac{\bm{v}}{\|\bm{v}\|}-\frac{\bm{v}_{\text{des}}}{\|\bm{v}_{\text{ref}}\|}\|$ 2.0-2.0- 2.0
@@ III-C2 Domain randomization
Mass [1.0,1.5][1.0,1.5][ 1.0 , 1.5 ] kg
Max. thrust [22,30][22,30][ 22 , 30 ] N
$I_{x},I_{y}$ inertial [0.075,0.125][0.075,0.125][ 0.075 , 0.125 ] $\text{kg}\cdot m^{2}$
$I_{z}$ inertial [0.15,0.25][0.15,0.25][ 0.15 , 0.25 ] $\text{kg}\cdot m^{2}$
Motor delay factor [0.5,0.8][0.5,0.8][ 0.5 , 0.8 ] -
Body rate delay factor [0.5,0.8][0.5,0.8][ 0.5 , 0.8 ] -
Drag force coefficient [0.4,0.6][0.4,0.6][ 0.4 , 0.6 ] -
@@ III-C3 Guiding drone to fly desired trajectories
Additionally, to ensure safe navigation, we incorporate an obstacle avoidance reward $r_{\text{obss}}$, which encourages the drone to maintain a safe distance from obstacles visible in its FOV. The reward formulation is as follows:
$r_{\text{obst}}=\bm{d}_{\text{obst}},\quad\bm{d}_{\text{obst}}<\bm{d}_{\text{th}},$ (17)
where $\bm{d}_{\text{obss}}$ represents the minimum distance to the nearest obstacle in drone’s FOV, $\bm{d}_{\text{th}}$ is the predefined threshold for proximity, which is set to 0.5m in our method.
@@ IV-A Training efficiency comparison with other methods
To validate GRaD-Nav’s training efficiency, we compared our method with other differentiable and non-differentiable RL algorithms. Back Propagation Through Time (BPTT) [37] is a critic-free differentiable reinforcement learning algorithm that passes gradients through the whole trajectory; Proximal Policy Optimization (PPO) [38] is a common non-differentiable, online reinforcement learning algorithm. All three algorithms were used for training policies that navigate through a trajectory of around 20m, passing through several challenging unstructured obstacles, which can be seen in 3. As much as possible, we used identical hyperparameters (Table VI) for different algorithms. It is to be noted that BPTT samples the whole trajectory for policy updating, meaning the horizon length equals the episode length, which can take up a lot of memory for each environment and heavily limits its maximum allowed number of parallel simulated environments. The experiment results in Fig.2 show that (i) non-differentiable RL can struggle to train a satisfactory policy for this end-to-end visual navigation task within $1\times 10^{7}$ samples; (ii) compared to vanilla BPTT, our method achieves over 300% sample efficiency and uses only 20% of training time to deliver a better policy.
@@ IV-B Ablation study of our methods
Trajectories Long traj. middle gate Long traj. Right gate
@@
The experiment results show that our proposed method achieves the highest training and evaluation rewards as well as success rate on both trajectories among all methods. As visual perception is our navigation policy’s major sensor input, it is not surprising that the policy without visual observation cannot conduct successful navigation. Using depth image rather than RGB image also demonstrated a low success rate in vision-ablated experiments, which emphasized the necessity of accessing RGB information for effective navigation on our task. As introduced in Section III-C3, drone’s body linear velocity plays an important role in our policy’s navigation-related reward items, velocity-ablated policies would suffer from serious compounding error, and thus cannot fly an ideal trajectory. Without CENet, our method can still train a policy network that achieves high rewards compared to other ablation cases. CENet ablated policies can complete a trajectory successfully when the initial condition is close to the ”neutral” initial condition (e.g. position = (0,0,1.3)m). However, policies without CENet can be vulnerable to any environmental perturbations, which seriously harm the robustness of the policy networks. All of the failure cases without CENet on two trajectories “crash” due to unsuccessful obstacle avoidance.
@@ IV-C Sim-to-real transfer of generalizable policy
Figure 4: A generalizable policy flying through gates positioned at varying locations with diverse distractor objects in the scene. The colored trajectory represents the drone’s velocity, with purple indicating low speed and yellow indicating high speed.
@@ V Conclusions
Limitations: Our method relies on hand-crafted reward shaping (e.g., trajectory waypoints), limiting it to single-task execution like gate traversal. It also depends on reliable velocity estimates from VIO, which may be inconsistent. Future work includes (i) multi-task training with language input, (ii) improving generalization via stronger backbones and diverse environments, and (iii) extending to contact-rich tasks such as mobile or aerial manipulation.
@@ VI Acknowledgment
The authors would like to thank Chenghao Zhu, Keiko Nagami and Javier Yu for fruitful discussion on FiGS. The GRaD-Nav algorithm was developed based on SHAC’s open-sourced codebase [30], $\beta$-VAE based CENet code was referred to [39].